OpenStudy (anonymous):

Mr. Doodle’s grade distribution over the past 3 years for a course in college algebra is shown in the chart below. Grade Number A 45 B 180 C 110 D 95 F 65 I 5 If Jane plans to take a college algebra course with Mr. Doodle, determine the empirical probability that she receives an A. please help and show work!

jimthompson5910 (jim_thompson5910):

There are 45 people who received A's There are a total of 45+180+110+95+65+5 = 500 grades So the probability is 45/500 = 9/100

OpenStudy (anonymous):

This question and answer was posted earlier...

OpenStudy (anonymous):

i wasnt sure if u were corrected!!!

OpenStudy (anonymous):

"people gave u same answer, how many do u want?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

look im sorry estudier i just want to make sure that it was correct okay!!

OpenStudy (anonymous):

Ok but u can imagine if everybody did that,,,

OpenStudy (anonymous):

true i agree that why i apologize !! i gave me the answer but i need the work as well!!!